shift summary, evening shift October 1

B. Cox (
Sat, 2 Oct 1999 00:27:34 -0500 (CDT)

Shift Summary for Evening Shift, Oct 1,/99

On Shift: Brad, Hogan

Hours of beam available at all intensities: 8.5
Hours of beam available at nominal intensity: 8.5
Hours of E832 data (special runs not included): 7.75
Hours of special runs (laser scans, muon runs...): 0.0

Accounting of lost time:

miscellaneous Kfrends/pipeline resets
change of tapes ~1/2 hour
processor accept fraction problem ~1/4 hour
Runs ended this shift:

Run 14788
Run 14789


Tapes taken to FCC this shift


Accesses during shift:


Overview of shift:

15:40 Called Main Control to ask that beam structure be worked on
Some improvement..

16:10 live time =68% at 8.8E12 Will keep monitoring it.

17:10 A number of kfrends required

19:20 improved spill structure leads to higher live time (65->70%%)
results in more events per unit beam leads to 1-2 seconds
trespassing on the next spill. Ask to lower beam intensity to
8.5x10**12 to compensate.

19:25 pipeline errors, reset pipeline and kfrend.
pipeline errors, reset pipeline and kfrend.

21:05 Changed tapes and started run 14789
Spill 1-5, run 14789, reset the HCC.
Had to do kfrends to get the run going, what else is new.

Adjusted the alarm limits for B01-B16/SEM. Tolerance = +/- 20%
about mean (was +/- 10% about mean ).

23:45 Noticed that there were problems with the L3 filter plane
accept fractions being different between processors. Do a complete
fresh start to try to clear up. Run 14879 will have to be inspected

Description of Shift:

Smooth until the very last accept fraction problem.
Still struggling at the end of shift

Follow up Notes on Previous Shift